Find the class boundaries, midpoints, and widths for each class.




1.58-62



2.125-131



3. 16.35-18.46



4. 16.3-18.5

1
Expert's answer
2021-12-20T19:00:41-0500

Solution:

1.

58-62

class boundaries: upper = 62 , lower = 58

midpoint = "\dfrac{62+58}2=60"

width = "62-58=4"


2.

125-131

class boundaries: upper =131, lower =125

midpoint ="\dfrac{131+125}2=128"

width ="131-125=6"


3.

16.35-18.46

class boundaries: upper = 18.46, lower =16.35

midpoint ="\dfrac{18.46+16.35}2=17.405"

width ="18.46-16.35=2.11"


4.

16.3-18.5

class boundaries: upper =18.5 , lower =16.3

midpoint = "\dfrac{18.5+16.3}2=17.4"

width = "18.5-16.3=2.2"